### v1

# TabNSA: Native Sparse Attention for Efficient Tabular Data Learning thanks: Code on GitHub: https://github.com/aseslamian/TabNSA

Abstract
Tabular data poses unique challenges for deep learning due to its heterogeneous features and lack of inherent spatial structure. This paper introduces TabNSA, a novel deep learning architecture leveraging Native Sparse Attention (NSA) specifically for efficient tabular data processing. TabNSA incorporates a dynamic hierarchical sparse strategy, combining coarse-grained feature compression with fine-grained feature selection to preserve both global context awareness and local precision. By dynamically focusing on relevant subsets of features, TabNSA effectively captures intricate feature interactions. Extensive experiments demonstrate that TabNSA consistently outperforms existing methods, including both deep learning architectures and ensemble decision trees, achieving state-of-the-art performance across various benchmark datasets.
Index Terms:
Tabular data, Native Sparse Attention
I Introduction
Tabular data, ubiquitous across diverse domains, often presents challenges for traditional deep learning models due to its heterogeneous nature and lack of inherent spatial structure. While deep learning has revolutionized fields like computer vision and natural language processing, its application to tabular data has been less successful, with ensemble decision trees continuing to dominate many real-world applications [1, 2]. This discrepancy stems from the fact that canonical architectures designed for encoding raw data into meaningful representations have not yet been effectively adapted for tabular data [3]. The relatively inferior performance of deep learning models on tabular data can be attributed to factors such as limited labeled training data and the models’ struggle to capture complex relationships [4]. Researchers are actively exploring innovative techniques to address the limitations of decision trees and expand the applicability of deep learning in this domain [5].
This paper introduces TabNSA, inspired by Natively Sparse Attention (NSA) [6], which adapts the sparse attention mechanism for tabular data. TabNSA employs dynamic sparse attention to process heterogeneous features efficiently.
Instead of treating all features uniformly, it learns to focus on the most relevant subset for each instance, dynamically adjusting feature importance. This hierarchical sparse attention allows TabNSA to capture both global context and fine-grained feature dependencies, performing coarse selection of the most informative features and then modeling their interactions in detail.
The key contributions of this work are as follows:
  • TabNSA Architecture: We propose TabNSA, a novel deep learning model that integrates Native Sparse Attention into tabular data learning. By dynamically selecting and attending to relevant subsets of features, TabNSA effectively addresses the challenge of feature heterogeneity and captures complex feature dependencies. This approach provides a strong foundation for multimodal applications.
  • State-of-the-Art Performance: Our comprehensive experiments demonstrate that TabNSA consistently outperforms existing methods, including both deep learning architectures and ensemble decision trees. TabNSA achieves state-of-the-art accuracy on most benchmark datasets, demonstrating its strong generalization capabilities across diverse tabular tasks.
II Related Work
Existing tabular learning methods aim to address various challenges associated with diverse datasets. These include situations where the feature sets differ between training and test data, limited or missing labels, and the potential for new features to emerge over time [7]. Classical techniques include parametric methods like Logistic Regression and non-parametric approaches such as K-Nearest Neighbors, Gradient Boosting, and Decision Trees [8]. Various models have been developed, including logistic regression, XGBoost [9, 10], and Multi-Layer Perceptrons.
Deep learning-based approaches have emerged to address the growing complexity and high-dimensionality of modern tabular datasets. These methods can automatically learn hierarchical representations and capture intricate patterns within the data [11]. Deep cross networks (DCN) [12] combine a deep network that learns high-order feature interactions with a cross-network that automatically applies feature crossing, efficiently capturing bounded-degree feature interactions. TabMixer [13] uses an enhanced MLP-Mixer architecture. Its design effectively extracts attention-aware features and captures complex interactions across tabular features and examples, even extending to dependencies between different tables.
Attention-based models, in particular, have gained popularity as they dynamically focus on the most relevant parts of the input data, although this can increase computational requirements. TabNet [3] is a attention-based mechanisms for tabular data processing. It concentrates on the most significant features at each decision stage, enabling efficient learning and providing interpretable feature attributions. Several models have been developed using Transformers as their foundational components. AutoInt [14] employs Transformers to determine the significance of diverse input features. TransTab [15] utilizes Transformers to handle tasks such as learning from multiple tables with partially overlapping columns. Similarly, TabTransformer [16] employs self-attention-based Transformers to enhance feature representations, surpassing classical methods in terms of accuracy, exhibiting robustness to missing data, and introducing a novel pre-training approach for semi-supervised learning. SAINT [17] uses a hybrid attention mechanism, focusing on both rows and columns with enhanced embeddings. It introduces a novel intersample attention approach and contrastive pre-training, outperforming gradient boosting methods.
III Background
To address the computational inefficiency of standard attention mechanisms in long-context modeling for tabular data processing, we adapt Native Sparse Attention—a natively trainable sparse attention architecture [18, 19]. This adaptation tackles the quadratic complexity of standard attention mechanisms by leveraging a hierarchical sparse strategy that combines token compression, blockwise token selection, and local sliding windows. Our goal is to enhance computational efficiency while maintaining both global and local context awareness in tabular data applications.
In the model, each sample’s feature vector is first embedded and then projected into queries, keys, and values:
$q=W_{q}X,\quad k=W_{k}X,\quad v=W_{v}X.$ (1)
This corresponds to the standard attention formulation. Here, $W_{q}$, $W_{k}$ and $W_{v}$ are learned projection matrices. Given an input sequence with query $q_{t}$, keys $K_{t}$, and values $v_{t}$ for $t$-features the full attention computes the output as:
$o_{t}$ $=\text{Attn}(q_{t},k_{1:t},v_{1:t})=\sum_{i=1}^{t}\alpha_{t,i}v_{i},$ (2)
$\text{with}\quad\alpha_{t,i}$ $=\frac{\exp\left(\frac{q_{t}k_{i}}{\sqrt{d_{k}}}\right)}{\sum_{j=1}^{t}\exp\left(\frac{q_{t}k_{j}}{\sqrt{d_{k}}}\right)}.$
NSA replaces the original key value pairs with a more compact representation $\tilde{k}_{t}\text{ and }\tilde{v}_{t}$ constructed as:
$\tilde{k}_{t}=f_{K}(q_{t},k_{1:t},v_{1:t}),\quad\tilde{v}_{t}=f_{V}(q_{t},k_{1:t},v_{1:t})$ (3)
The functions $f_{K}$ and $f_{V}$ implement three distinct remapping strategies $\mathcal{C}=\{cmp,slc,win\}$ representing compression, selection, and sliding window, respectively:
  1. 1.
    Token Compression: Aggregates sequential blocks of keys or values into block-level representations. The compressed key representation is defined as:
    $\tilde{k}_{t}^{\text{cmp}}=f_{K}^{\text{cmp}}(k_{1:t})=\left\{\varphi\left(k_{id+1:id+l}\right)\middle|0\leq i\leq\left\lfloor\frac{t-L}{d}\right\rfloor\right\}$ (4)
    where $l$ is the block length, $d$ is the sliding stride between adjacent blocks, and $\varphi$ is a learnable MLP with intra-block position encoding.
  2. 2.
    Token Selection: Selectively preserves individual keys and values. The selection strategy processes key and value sequences in spatial continuous blocks. The importance scores for selection blocks are derived from the attention computation of compression tokens:
    $\mathbf{p}_{t}^{\text{cmp}}=\text{Softmax}\left(\mathbf{q}_{t}^{T}\tilde{k}_{t}^{\text{cmp}}\right)$ (5)
    $\mathbf{p}_{t}^{\text{cmp}}$ denote the attention scores between $q_{t}$ and the compressed keys.
    Let $l^{\prime}$ denote the selection block size. By using the same blocking pattern for both compression and selection blocks, we can directly obtain the importance scores for the selection blocks $\mathbf{p}_{t}^{\text{slc}}=\mathbf{p}_{t}^{\text{cmp}}$. When the blocking schemes vary, we derive the importance scores for selection blocks based on their spatial relationships.So, the selection score for block $j$ can be computed as:
    $\mathbf{p}_{t}^{\text{slc}}[j]=\sum_{m=0}^{\frac{l^{\prime}}{d}-1}\sum_{n=0}^{\frac{l}{d}-1}\mathbf{p}_{t}^{\text{cmp}}\left[\frac{l^{\prime}}{d}j-m-n\right]$ (6)
[…]
    $\mathcal{I}_{t}=\{i\mid\operatorname{rank}(\mathbf{p}_{t}^{\text{slc}^{\prime}}[i])\leq n\}$ (7)
    $\tilde{k}_{t}^{\text{slc}}=\operatorname{Cat}\left[\{k_{il^{\prime}+1:(i+1)l^{\prime}}\mid i\in\mathcal{I}_{t}\}\right],$ (8)
[…]
  3. 3.
    Sliding Window: To handle local patterns effectively, a fixed-size sliding window of recent tokens is maintained:
    $\tilde{k}_{t}^{\text{win}}=k_{t-w:t},\quad\tilde{v}_{t}^{\text{win}}=v_{t-w:t}$ (9)
    After the sparse attention module combines the outputs from the different branches, the resulting representation is processed by further linear layers (and activation functions like GELU) to produce the final regression output. The final output is as follows:
    $o_{t}^{*}=\sum_{c\in\mathcal{C}}g_{t}^{c}\cdot\text{Attn}(q_{t},\tilde{K}_{t}^{c},\tilde{V}_{t}^{c})$ (10)
    $\hat{y}=f(o_{t}^{*}).$ (11)
    The gate score, $g^{c}_{t}\in[0,1]$, is obtained from the input features using sigmoid activation and an MLP and $f$ is final linear layer.
For tabular data learning, NSA can be adapted by treating each feature as a token and learning the relationships between these features through its sparse attention mechanisms. This approach is particularly beneficial for high-dimensional tabular data, where the computationally intensive standard attention mechanisms would be prohibitive.
IV Experiments and Results
IV-A Dataset
In this work, we evaluated the performance of the TabNSA model and several baseline models across a diverse set of datasets. The assessment included 8 widely used binary classification datasets, 2 multi-class datasets, and 3 regression datasets sourced from the UCI repository, OpenML, and Kaggle. These datasets span a wide range of domains, including finance, business, chemistry, geography, ecology, image recognition, and sports, ensuring a comprehensive evaluation of the models. This diverse benchmark dataset collection enables us to systematically assess the performance across various domains and data structures. A detailed description of each dataset is provided in Table I.
TABLE I: Dataset details including abbreviation, number of classes, number of data points, and number of features.
Dataset Name Abbreviation # Class # Data # Features
Credit-g CG 2 1,000 20
Credit-Approval CA 2 690 15
Dataset-Sales DS 2 690 15
Adult AD 2 48,842 14
Cylinder-Bands CB 2 540 35
Blastchar BL 2 7,043 35
Insurance-Co IO 2 5,822 85
1995-Income IC 2 32,561 14
ImageSegmentation SG 7 2,310 20
ForestCovertype FO 7 581,012 55
Sarcos-Robotics SA - 48,933 28
Topo-2-1 TO - 8,885 266
CPU-Predict CP - 8,192 13
IV-B Preprocessing
In order to handle missing values and class imbalance we used the preprocessing method in [13]. Let the input variable space as $\mathcal{D}ata\in\{\mathbb{R}\cup\mathbb{C}\cup\mathbb{B}\cup\varnothing\}$, where $\mathbb{R}$, $\mathbb{C}$, $\mathbb{B}$ and $\varnothing$ represent the domains of numerical, categorical, binary dat and NULL values respectively. After preprocessing block, we have $\{\mathcal{X},\mathcal{Y}\}$ denote a feature-target pair, where $x$ represents numerical features, and $y$ is a continuous variable for regression tasks or a whole number for classification tasks.
IV-C Experiment Set up
To enhance TabNSA’s performance and stability, we systematically tuned critical hyperparameters using OPTUNA [20]. This involved dividing each dataset into training (70%), validation (10%), and test (20%) sets, training TabNSA on the training set, and evaluating it on the validation set to guide the hyperparameter optimization process.
We employed the OPTUNA package to optimize key parameters, ensuring an efficient and automated search for the best configuration. Hyperparameter optimization is crucial for tabular data, given the variability in feature distributions and dataset characteristics. Once the optimal parameters are determined, the final model is tuned on train dataset and evaluated on the test set using standard metrics to assess its generalization performance. The search space included architectural parameters such as head dimension (8-46), number of attention heads (1-8), sliding window size (1-8), compression block size (4-16), selection block size (2-compression block size), and number of selected blocks (1-4). Furthermore, we fine-tuned optimization parameters like learning rate (1e-4 - 1e-3) and batch size (32-128) to strike a balance between convergence speed and generalization.
IV-D Binary Classification
We evaluated TabNSA across eight datasets for the supervised binary classification task, with the results summarized in Table II. We include the results of other methods directly from [15]. To ensure a fair comparison, we followed the instructions in [15], ran the model 10 times, and reported the average. The model consistently outperformed all competing approaches, including classical machine learning methods, deep neural networks (DNNs), transformer-based architectures, and tree-based models.
Notably, TabNSA ranked first across all datasets, demonstrating its effectiveness. While it may not achieve the highest performance on every dataset, it consistently excels overall.
TABLE II: Performance Comparison of Models for Binary Classification
Methods CG CA DS AD CB BL IO IC Rank (Std) $\downarrow$
Logistic Regression 0.720 0.836 0.557 0.851 0.748 0.801 0.769 0.860 11.5 (2.33)
XGBoost 0.726 0.895 0.587 0.912 0.892 0.821 0.758 0.925 6.69 (4.05)
MLP 0.643 0.832 0.568 0.904 0.613 0.832 0.779 0.893 10.9 (2.01)
SNN 0.641 0.880 0.540 0.902 0.621 0.834 0.794 0.892 9.81 (3.66)
TabNet 0.585 0.800 0.478 0.904 0.680 0.819 0.742 0.896 12.7 (1.67)
DCN 0.739 0.870 0.674 0.913 0.848 0.840 0.768 0.915 5.5 (2.41)
AutoInt 0.744 0.866 0.672 0.913 0.808 0.844 0.762 0.916 5.5 (2.83)
TabTrans 0.718 0.860 0.648 0.914 0.855 0.820 0.794 0.882 7.56 (4.12)
FT-Trans 0.739 0.859 0.657 0.913 0.862 0.841 0.793 0.915 5.5 (2.46)
VIME 0.735 0.852 0.485 0.912 0.769 0.837 0.786 0.908 8.31 (2.55)
SCARF 0.733 0.861 0.663 0.911 0.719 0.833 0.758 0.919 7.88 (2.89)
TransTab 0.768 0.881 0.643 0.907 0.851 0.845 0.822 0.919 4.56 (2.61)
TabMixer 0.660 0.907 0.659 0.900 0.829 0.821 0.974 0.969 6.19 (4.80)
TabNSA 0.874 0.898 0.717 0.967 0.721 0.926 0.979 0.967 2.38 (3.11)
IV-E Multi-class Classification
Table III presents a performance comparison of TabNSA and other neural network models on two multi-class classification datasets. Since multi-class classification tasks often suffer from class imbalance, we used macro-F1 as the primary evaluation metric during training to ensure a balanced assessment across all classes, following [21]. The results show that TabNSA consistently outperforms all models on both datasets.
TABLE III: Comparison of different methods on SG and FO datasets.
Dataset Metrics Method
MLP TabTrans TabNet TabNSA
SG ACC 90.97 - 96.09 98.27
F1 90.73 - 94.96 98.25
FO ACC 67.09 68.76 65.09 92.89
F1 48.03 49.47 52.52 82.66
IV-F Regression Task
To evaluate model performance on a regression task, we compared five methods—MLP, XGBoost, CatBoost, TabNet, and our proposed TabNSA—using Root Mean Squared Error (RMSE) as the metric across three datasets: SA, TO, and CP. Lower RMSE indicates better performance. As shown in Table IV, XGBoost achieved the best overall performance, with the lowest RMSE on the CP dataset and competitive scores across the others, resulting in the highest ranking. CatBoost closely followed, outperforming XGBoost on the SA dataset and ranking second overall. The MLP model struggled with relatively higher RMSE values, while TabNet exhibited significantly worse performance, particularly on the SA dataset. Our TabNSA model achieved the lowest RMSE on the TO dataset, demonstrating its effectiveness in certain cases. However, it reported a higher RMSE on the CP dataset, impacting its overall ranking. These results highlight the effectiveness of gradient-boosting models for tabular regression, while also showcasing TabNSA’s strengths in specific domains.
TABLE IV: Evaluation of Different Models for Regression task. Results shows RMSE
Methods SA TO CP Rank $\downarrow$
MLP 0.607 0.037 7.895 4
XGBoost 0.185 0.0276 2.631 1.67
CatBoost 0.183 0.0281 2.848 2
TabNet 3.2450 0.1050 3.405 4.33
TabNSA 0.3129 0.0273 14.2931 3
V Conclusion
In this paper, we introduced TabNSA, a novel deep learning framework that adapts Native Sparse Attention (NSA) for tabular data learning. TabNSA addresses the computational inefficiencies of traditional attention mechanisms by dynamically focusing on relevant features through a hierarchical sparse strategy combining token compression, selection, and sliding windows.
Our experiments demonstrate TabNSA’s superior performance on classification tasks, while showing promising results on certain regression datasets that suggest potential for further improvement. TabNSA’s advantages over previous Transformer-based approaches include reduced computational complexity, hardware-aligned design, and native integration of sparsity during training. These properties enable the model to effectively balance global context with local feature interactions.
TabNSA features reduced computational complexity, hardware-aligned design, and inherent sparsity during training, allowing it to effectively balance global context and local feature interactions.